### v1

# SPLITZ: Certifiable Robustness via Split Lipschitz Randomized Smoothing thanks: This work was supported by NSF grants CCF 2100013, CNS 2209951, CCF 1651492, CNS 2317192 and CNS 1822071. This wrok has been submitted to IEEE Transactions on Information Forensics and Security

Abstract
Certifiable robustness gives the guarantee that small perturbations around an input to a classifier will not change the prediction. There are two approaches to provide certifiable robustness to adversarial examples– a) explicitly training classifiers with small Lipschitz constants, and b) Randomized smoothing, which adds random noise to the input to create a smooth classifier. We propose SPLITZ, a practical and novel approach which leverages the synergistic benefits of both the above ideas into a single framework. Our main idea is to split a classifier into two halves, constrain the Lipschitz constant of the first half, and smooth the second half via randomization. Motivation for SPLITZ comes from the observation that many standard deep networks exhibit heterogeneity in Lipschitz constants across layers. SPLITZ can exploit this heterogeneity while inheriting the scalability of randomized smoothing. We present a principled approach to train SPLITZ and provide theoretical analysis to derive certified robustness guarantees during inference. We present a comprehensive comparison of robustness-accuracy tradeoffs and show that SPLITZ consistently improves upon existing state-of-the-art approaches on MNIST and CIFAR-10 datasets. For instance, with $\ell_{2}$ norm perturbation budget of $\epsilon=1$, SPLITZ achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%.
Index Terms:
Certified defense, Randomized smoothing, Lipschitz constants, Adversarial defense.
I Introduction
As deep learning becomes dominant in many important areas, ensuring robustness during test time becomes increasingly important. Deep neural networks are known to be vulnerable to adversarial attacks: small imperceptible perturbations in the inputs leading to incorrect decisions [1, 2]. Although many works have proposed heuristic defenses for training robust classifiers, they are often shown to be inadequate against adaptive attacks [3, 4, 5]. Therefore, a growing literature on certifiable robustness has emerged [6, 7, 8]; where the classifier’s prediction must be provably robust around any input within a perturbation budget.
There are two broad approaches to design classifiers which are certifiably robust: a) design classifiers which are inherently stable (i.e., smaller Lipschitz constants) [9, 10, 11]. There are a variety of methods to train classifiers while keeping the Lipschitz constants bounded. The second approach is b) randomized smoothing (RS) [7, 12, 8]; here, the idea is to smooth the decision of a base classifier by adding noise at the input. The approach of RS has been generalized in several directions: Salman et al. [13] and Carlini et al. [14] combine denoising mechanisms with smoothed classifiers, Salman et al. [15] combine adversarial training with smoothed classifiers, Zhai et al. [16] propose a regularization which maximizes the approximate certified radius and Horváth et al. [17] combine ensemble models with smoothed classifiers.
Ensuring certified robustness by constraining the Lipschitz constant usually involves the estimation of the Lipschitz constant for an arbitrary neural network. The main challenge is that accurate estimation of Lipschitz constants becomes hard for larger networks, and upper bounds become loose leading to vacuous bounds on certified radius. Recently, a variety of approaches have emerged that focus on training while explicitly constraining the model’s Lipschitz constant (outputs), for instance, LMT [18] and BCP [11]. These methods control the Lischitz constant of the model by constraining the outputs of each layer (or the outputs of the model); this has the dual benefit of enhanced robustness of the model as well as better estimation of the overall Lipschitz constant of the model. To further minimize the Lipschitz constant during the certified robust training process and better estimate the local Lipschitz constant of the model, recent work [19] focus on the constrained training with respect to the local Lipschitz constant by utilizing the clipped version of the activation functions.
Another line of works [20, 21] propose using models for which each individual layer is $1$-Lipschitz. By enforcing orthogonal or near-orthogonal weight matrices, these networks naturally limit their sensitivity to input perturbations, contributing to a form of robustness that does not solely depend on Lipschitz constant estimation. This approach can complement the limitations of both Lipschitz constrained training and RS, particularly in handling larger models where direct Lipschitz estimation becomes impractical. Orthogonal constraints help maintain the global Lipschitz constant close to $1$.
Lipschitz constrained training provides deterministic guarantees on certified radius and is often challenging to accurately estimate the Lipschitz constant of a large neural network. RS on the other hand offers scalability to arbitrarily large networks and provide the closed-form certified robust radius. These guarantees, however, are probabilistic in nature and the smoothing procedure treats the entire classifier as a black box mechanism.
TABLE I: Comparison of certified test accuracy (%) on MNIST under $\ell_{2}$ norm perturbation. Each entry lists the certified accuracy using numbers taken from respective papers (RS results follow from previous benchmark papers [22, 12]).
Certified Test Accuracy at $\epsilon~{}(\%)$
Method Extra data 1.50 1.75 2.00 2.25 2.50
RS [7] 67.3 46.2 32.5 19.7 10.9
MACER [16] 73.0 50.0 36.0 28.0 -
Consistency [22] 82.2 70.5 45.5 37.2 28.0
SmoothMix [12] 81.8 70.7 44.9 37.1 29.3
DRT [23] 83.3 69.6 48.3 40.3 34.8
RS+OrthoNN [21] 70.1 49.7 33.2 21.0 10.5
SPLITZ 80.2 71.3 62.3 51.7 40.5
Figure 1: (a) Lipschitz constants of each affine layer in pretrained models: VGG16 [24], ResNet18 [25], SqueezeNet [26]. (b) Local Lipschitz (upper) bound for three random CIFAR-10 images on VGG16; (c) Percentage analysis of local Lipschitz (upper) bound in CIFAR-10 test data (additional results are presented in the Appendix A).
Overview of SPLITZ and Contributions. In this paper, we propose SPLITZ, which combines and leverages the synergies offered by both local Lipschitz constrained training and randomized smoothing. The general idea is to split a classifier into two halves: the first half (usually a few layers) is constrained to keep a smaller Lipschitz (upper) bound, and the latter half of the network is smoothed via randomization. We propose the use of local Lipschitz constant(s) of the first half of the network. This is because it can capture the stability of the model with respect to each individual input. To enhance certified robustness, we incorporated the local Lipschitz constant of the network’s first half into the loss function as a regularization term. This approach aims to maintain a comparatively small local Lipschitz constant (typically less than 1) for the network’s first half, thereby improving the certified robustness. To obtain the certified radius during inference, we derive a closed-form expression for the certified radius based on the local Lipschitz constant as well as the randomized smoothing parameters as shown in Theorem 1; this result illustrates that enforcing a relatively small local Lipschitz constant can help in improving the certified radius. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing.
Interestingly, this approach yields state-of-the-art results for several datasets. For instance, Table I compares the certified test accuracy of SPLITZ and existing state-of-art techniques for various values of $\epsilon$ (perturbation budget or certified radius) on the MNIST dataset. For $\epsilon$ ($\ell_{2}$ norm perturbation size) even as large as $2.5$, where the state-of-the-art accuracy is $34.8\%$, SPLITZ achieves certified accuracy of around $40.5\%$. In Section IV, we present comprehensive set of results on MNIST and CIFAR-10 datasets. In addition, we also provide a detailed ablation study, and study the impact of various hyperparameters (such as the location of the split, randomized smoothing parameters).
Intuition behind SPLITZ. The intuition behind SPLITZ comes from the following key observations: a) Layer-wise Heterogeneity: many modern deep networks exhibit heterogeneity in Lipschitz constants across layers. Fig 1(a) shows the per-layer Lipschitz constants for three networks (VGG16, ResNet18 and SqueezeNet). We observe that the values can vary widely across the layers, and quite often, latter half of the network often shows larger Lipschitz constants. b) Input (local) heterogeneity: We show the local Lipschitz (upper) bounds for three randomly sampled images from CIFAR-10 when passed through the first four layers of VGG16; note that the values of local Lipschitz bound can vary across different inputs (images). The same behavior across the entire CIFAR-10 test dataset is shown in Fig. 1(c). These observations motivate SPLITZ as follows: smoothing the input directly may not be the optimal approach as it does not account for this heterogeneity. Instead, by introducing noise at an intermediate stage of the classifier, the model can become more resilient to perturbations. This suggests the idea of splitting the classifier. Simultaneously, the first half of the network should also be “stable”, which motivates constraining the Lipschitz bound of first half of the network.
The paper is organized as follows: Section II introduces the objectives of the paper, preliminaries, and definitions of randomized smoothing and Lipschitz constant(s) used in the paper. Additionally, we review related works concerning randomized smoothing and Lipschitz training. Section III discusses the theoretical guarantees of SPLITZ and the corresponding training mechanisms. Experiments and evaluation results are presented and discussed in Section IV. Furthermore, additional experimental details on the Lipschitz constant, theoretical results on the local Lipschitz bound, and supplementary experimental results of the main findings are provided in the Appendices A, B, and C, respectively. Our code is available at: https://anonymous.4open.science/r/SPLITZ-3C66.
II Preliminaries on Certified Robustness
We consider a robust training problem for multi-class supervised classification, where we are given a dataset of size $N$, $\{x_{i},y_{i}\}_{i=1}^{N}$, where $x_{i}\in\mathbb{R}^{d}$ denotes the set of features of the $i$th training sample, and $y_{i}\in\mathcal{Y}:=\{1,2,\dots,C\}$ represents the corresponding true label. We use $f$ to denote a classifier, which is a mapping $f$: $\mathbb{R}^{d}\rightarrow\mathcal{Y}$ from input data space to output labels. From the scope of this paper, our goal is to learn a classifier which satisfies certified robustness, as defined next.
Definition 1.
(Certified Robustness) A (randomized) classifier $f$ satisfies $(\epsilon,\alpha)$ certified robustness if for any input $x$, we have
$\mathbb{P}(f(x)=f(x^{\prime}))\geq 1-\alpha$, $\forall x^{\prime}$, such that $x^{\prime}=x+\delta,\parallel\delta\parallel_{p}\leq\epsilon$
where the probability above is computed w.r.t. randomness of the classifier $f$.
Intuitively, certified robustness requires that for any test input $x$, the classifier’s decision remains locally invariant, i.e., for all $\forall x^{\prime}$ around $x$, such that $\parallel x^{\prime}-x\parallel_{p}\leq\epsilon$, $f(x)=f(x^{\prime})$ with a high probability. Thus, $\epsilon$ is referred to as the certified radius, and $(1-\alpha)$ measures the confidence. We mainly focus on $\ell_{2}$ norm ($p=2$) for the scope of this paper.
The literature on certified robustness has largely evolved around two distinct techniques: Randomized Smoothing and Lipschitz constrained training for Certifiably Robustness. We first briefly summarize and give an overview of these two frameworks, before presenting our proposed approach of Split Lipschitz Smoothing.
Randomized Smoothing (RS) [7] is a general procedure, which takes an arbitrary classifier (base classifier) $f$, and converts it into a ”smooth” version classifier (smooth classifier). Most importantly, the smooth classifier preserves nice certified robustness property and provides easily computed closed-form certified radius. Specifically, a general smooth classifier $g_{RS}(\cdot)$ derived from $f$ is given as:
$g_{RS}(x)=\underset{c~{}\in~{}\mathcal{Y}}{\text{argmax}}~{}\underset{\delta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{P}}(f(x+\delta)=c)$ (1)
Intuitively, for an input $x$, $g(x)$ will output the most probable class predicted by the base classifier $f$ in the neighbourhood of $x$ with a high confidence $1-\alpha$. In the paper [7], they prove that $g(x)$ is robust against $\ell_{2}$ perturbation ball of radius $\epsilon=\sigma\Phi^{-1}(\underline{p_{A}})$ around $x$, where $\sigma$ is the standard deviation of the Gaussian noise, and $\underline{p_{A}}$ is the lower bound of the probability that the most probable class predicted by the classifier $f$ is $c_{A}$. RS is arguably the only certified defense which can scale to large image classification datasets. Based on RS, a number of studies have been undertaken in this field: RS was originally proposed to deal with $\ell_{2}$ norm bounded perturbations; but was subsequently extended to other norms using different smoothing distributions, including $\ell_{0}$ norm with a discrete distribution [27], $\ell_{1}$ norm with a Laplace distribution [28], and the $\ell_{\infty}$ norm with a generalized Gaussian distribution [29]. Other generalizations include combining RS with adversarial training to further improve certified robustness and generalization performance [15] or denoising mechanisms (such as diffusion models) are often considered in conjunction with RS [13, 14].
Achieving a large certified radius can be equivalently viewed as learning a classifier with small Lipschitz constant. The Lipschitz constant is a fundamental factor in numerous studies focused on training a certifiably robust neural network, which can be defined as follows:
Definition 2.
(Global and Local Lipschitz Constant(s)) For a function $f$: $\mathbb{R}^{d}\rightarrow\mathcal{Y}$, the Global, Local, and $\gamma$-Local Lipschitz constants (at an input $x$) are respectively, defined as follows:
(Global Lipschitz constant)
$~{}~{}~{}~{}~{}~{}L_{f}=\underset{x,y\in\text{dom}(f);x\neq y}{\sup}\frac{||f(y)-f(x)||_{p}}{||y-x||_{p}}$ (2)
(Local Lipschitz constant)
$~{}~{}~{}~{}~{}~{}L_{f}(x)=\underset{y\in\text{dom}(f);y\neq x}{\sup}\frac{||f(y)-f(x)||_{p}}{||y-x||_{p}}$ (3)
($\gamma$-Local Lipschitz constant)
$~{}~{}~{}~{}~{}~{}L^{(\gamma)}_{f}(x)=\underset{y\in B(x,\gamma);y\neq x}{\sup}\frac{||f(y)-f(x)||_{p}}{||y-x||_{p}},$ (4)
where $B(x,\gamma)$ denotes the $\ell_{p}$-ball around $x$ of radius $\gamma$, i.e., $B(x,\gamma)=\{u:||u-x||_{p}\leq\gamma\}$.
Informally, $L^{(\gamma)}_{f}(x)$ captures the stability of the function $f$ in the neighborhood of $x$, where the neighborhood is characterized by an $\ell_{p}$-ball of radius $\gamma$.
Figure 2: Schematic of SPLITZ training framework. We first feed the input $x$ to the left half of the classifier denoted as $f_{L}$, where the local Lipschitz constant of $f_{L}$ is constrained. Subsequently, we smooth the right half of the classifier by introducing noise to the output of the left half, expressed as: $f_{L}(x)+\delta$. Finally, the output of the right half is $f_{R}(f_{L}(x)+\delta)$.
Lipschitz constrained training for Certifiably Robustness A reliable upper bound for the local Lipschitz constant is essential for the robustness of a classifier. However, computing the exact value of local Lipschitz constants can be computationally challenging, prompting researchers to seek approximations, in terms of upper bounds. Thus, a line of works focus on deriving a tighter local Lipschitz bound e.g., [30, 31, 32]. Another line of works utilize the local Lipschitz bound to obtain better robustness guarantees, e.g., [33, 34]. Furthermore, there are several works which aim to train a certified robust classifier as we briefly summarize next. One approach is to estimate/upper bound the global Lipschitz constant of the classifier (during each training epoch) and use it to ensure robustness. For instance, [18, 11, 35] follow this general approach. The challenge is that the bounds on global Lipschitz constants can be quite large, and do not necessarily translate to improve certified robustness. An alternative approach is to use a local Lipschitz bound (for each individual input $x$), as in [19] and then explicitly minimize the Lipschitz bound during the training process. For simplicity, we refer to the upper bound of the local Lipschitz constant as the “local Lipschitz constant”.
III SPLITZ: Inference, Certification and Training
In this Section, we first describe the details of the proposed SPLITZ classifier along with the motivation as well as key distinctions from prior work. We then present new theoretical results on certified radius for SPLITZ. Subsequently, we describe the training methodology for SPLITZ as well as inference and computation of the certified radius. Suppose we are given a base classifier $f:\mathbb{R}^{d}\rightarrow\mathcal{Y}$ which is a composition of $K$ functions. Consider an arbitrary “split” of $f$ as $f(\cdot)=f_{\text{R}}(f_{\text{L}}(\cdot))\triangleq f_{\text{R}}\circ f_{\text{L}}$. As an example, if the classifier has $K=2$ hidden layers, i.e., $f(x)=f_{2}(f_{1}(x))$, then there are $K+1=3$ possible compositions/splits:
$a)f_{\text{R}}=I,f_{\text{L}}=f_{2}\circ f_{1},$
$b)f_{\text{R}}=f_{2},f_{\text{L}}=f_{1},$
$c)f_{\text{R}}=f_{2}\circ f_{1},f_{\text{L}}=I,$
where $I$ represents the identity function.
Definition 3.
(SPLITZ Classifier) Let $f$ be a base classifier: $\mathbb{R}^{d}\rightarrow\mathcal{Y}$. Consider an arbitrary split of $f$ as $f(\cdot)=f_{\text{R}}(f_{\text{L}}(\cdot))$. We define the SPLITZ classifier $g_{\textit{SPLITZ}}(\cdot)$ as follows:
$g_{\textit{SPLITZ}}(x)=\underset{c~{}\in~{}\mathcal{Y}}{argmax}~{}\underset{\delta\sim\mathcal{N}(0,\sigma^{2}I)}{\mathbb{P}}(f_{\text{R}}(f_{\text{L}}(x)+\delta)=c)$ (5)
The SPLITZ smoothing classifier is illustrated in Fig 2. The basic idea of SPLITZ is two fold: smooth the right half of the network using randomized smoothing and constrain the Lipschitz constant of the left half. Specifically, to robustly classify an input $x$, we add noise to the output of the left half (equivalently, input to the right half) of the network, i.e., $f_{\text{L}}(x)$ and then follow the same strategy as randomized smoothing thereafter. While RS takes care of smoothing the right half, we would still like the left half to be as stable as possible. Thus in addition to smoothing, we need to ensure that the Lipschitz constant of the left half $f_{\text{L}}$ of the network is also kept small. We next present our main theoretical result, which allows us to compute the certified radius for SPLITZ.
Theorem 1.
Let us denote $L_{f_{\text{L}}}^{(\gamma)}(x)$ as the $\gamma$-local Lipschitz constant of the function $f_{\text{L}}$ at $x$ in a ball of size $\gamma$, and $R_{f_{\text{R}}}(f_{\text{L}}(x))$ as the certified radius of the function $f_{\text{R}}$ at the input $f_{\text{L}}(x)$, with probability at least $(1-\alpha)$. Then, for any input $x$, with probability $1-\alpha$, $g_{\textit{SPLITZ}}(x)$ has a certified radius of at least,
$R_{g_{\textit{SPLITZ}}}(x)=\underset{\gamma\geq 0}{\max}~{}~{}\min\left\{\frac{R_{f_{\text{R}}}(f_{\text{L}}(x))}{L_{f_{\text{L}}}^{(\gamma)}(x)},\gamma\right\}$ (6)
Figure 3: Comparison of certified radius with $\ell_{2}$ norm perturbation w.r.t RS [7] and SPLITZ (ours), when varying levels of Gaussian noise $\sigma\in\{0.25,0.5,1.0\}$ on the CIFAR-10 dataset. We can observe that SPLITZ consistently outperforms RS [7] under different noise levels. For instance, when Radius = 1.0, SPLITZ achieves $43.2\%$ certified test accuracy while the best certified test accuracy of RS under different noise levels is $22.0\%$. We refer the reader to Table II for a comprehensive comparison of SPLITZ with several other approaches on CIFAR-10 dataset.
Given an input $x$, in order to compute the certified radius for SPLITZ classifier, we need $L_{f_{\text{L}}}^{(\gamma)}(x)$, i.e., the $\gamma$-local Lipschitz constant (discussed in the next Section) and the certified radius of right half of the classifier, i.e., $R_{f_{\text{R}}}(f_{\text{L}}(x))$. For Gaussian noise perturbation in the second half, $R_{f_{\text{R}}}(f_{\text{L}}(x))$ is exactly the randomized smoothing $\ell_{2}$ radius [7], given as $R_{f_{\text{R}}}(f_{\text{L}}(x))=\frac{\sigma}{2}(\Phi^{-1}(\underline{p_{A}})-\Phi^{-1}(\overline{p_{B}}))$, where $\underline{p_{A}}$ is the lower bound of the probability that the most probable class $c_{A}$ is returned, $\underline{p_{B}}$ is upper bound of the probability that the “runner-up” class $c_{B}$ is returned.
Remark 1.
Optimization over $\gamma$ We note from Theorem 1 that finding the optimal choice of $\gamma$ is crucial. One way is to apply the efficient binary search during the certify process to find the optimal value of $\gamma$. Specifically, we set the initial value of $\gamma$ and compute the corresponding local Lipschitz constant $L_{f_{L}}^{(\gamma)}(x)$ at input $x$. By comparing the value between $\gamma$ and ${R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma)}(x)}$, we divide the search space into two halves at each iteration to narrow down the search space until $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{L_{f_{L}}^{(\gamma^{*})}(x)}$. Another way is to do a one-step search. Specifically, we first approximate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma)}(x)$ at $x$ by averaging local Lipschitz constants of inference data given the inference $\gamma$. We then set $\gamma^{\prime}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma)}(x)}$ and re-calculate the local Lipschitz constant $\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)$ according to $\gamma^{\prime}$. Finally, we compute the approximate optimal $\gamma^{*}={R_{f_{R}}(f_{L}(x))}/{\tilde{L}_{f_{L}}^{(\gamma^{\prime})}(x)}$. Overall, we show the certification process in detail in Algorithm 1.
Remark 2.
[…]
Remark 3.
Compatibility of SPLITZ with other defenses In addition, we claim that the SPLITZ mechanism is also compatible with other RS based certified robust techniques, such as adversarial smoothing [15], mixsmoothing [12] or denoising diffusion models [14]. As an example, [14] propose a denosing mechanism using a diffusion model, which achieves the state-of-the-art. Our SPLITZ classifier contains two parts, left part is constrained by a small local Lipschitz constant while right part is smoothed by noise, which is same as a randomized smoothing based mechanism. Thus, our model can easily add a diffusion denoising model after the noise layer (after $f_{L}(x)+\delta$) and then feed the denoised samples into the $f_{R}$. Similarly, for adversarial smoothing or mixsmoothing, SPLITZ is adaptable to feed either adversarial examples ($f_{L}(x^{\prime})+\delta$) or mixup samples ($f_{L}(\tilde{x})+\delta$) respectively to the right half of the classifier $f_{R}$. On the other hand, SPLITZ is also compatible to other Lipschitz constrained based mechanisms (e.g., orthogonal neural network based mechanism [21, 20]). For instance, we can incorporate an orthogonal constraint regularization into the loss function to enforce orthogonality within the convolution layer throughout the training process. The outcomes of integrating the orthogonal convolution neural network with SPLITZ are meticulously detailed in the Section IV.
We next present our proof of Theorem 1 as follows.
Proof.
Let us consider an input $x$ to SPLITZ classifier $g_{\textit{SPLITZ}}(\cdot)$ and define the following function
$\tilde{g}(u)\triangleq\underset{c\in\mathcal{Y}}{argmax}~{}\mathbb{P}_{\delta}(f_{R}(u+\delta)=c).$ (7)
We first note from Definition 3 that $g_{\textit{SPLITZ}}(x)$ can be written as $g_{\textit{SPLITZ}}(x)=\tilde{g}(f_{L}(x))$, where the function $\tilde{g}$ is the smoothed version of $f_{R}$.
We are given that the smooth version $\tilde{g}$ has a certified radius of $R_{\tilde{g}}(u)\triangleq R_{f_{\text{R}}}(u)$ with probability at least $1-\alpha$. This is equivalent to the statement that for all $u^{\prime}$ such that $||u-u^{\prime}||_{p}\leq R_{f_{\text{R}}}(u)$, we have:
$\tilde{g}(u)=\tilde{g}(u^{\prime}).$ (8)
We are also given $L_{f_{\text{L}}}^{(\gamma)}(x)$, the $\gamma$-local Lipschitz constant of the function $f_{\text{L}}$ at $x$ in a ball of size $\gamma$. This implies that for all $||x-x^{\prime}||_{p}\leq\gamma$,
$||f_{L}(x)-f_{L}(x^{\prime})||_{p}\leq L_{f_{L}}^{(\gamma)}(x)||x-x^{\prime}||_{p}$ (9)
Now, setting $u=f_{\text{L}}(x)$ and $u^{\prime}=f_{\text{L}}(x^{\prime})$, we obtain
$||u-u^{\prime}||_{p}=||f_{L}(x)-f_{L}(x^{\prime})||_{p}\leq L_{f_{L}}^{(\gamma)}(x)||x-x^{\prime}||_{p}.$ (10)
Now observe that ensuring $g_{SPLITZ}(x)=g_{SPLITZ}(x^{\prime})$ is equivalent to ensuring $\tilde{g}(f_{L}(x))=\tilde{g}(f_{L}(x^{\prime}))$, which in turn is equivalent to $\tilde{g}(u)=\tilde{g}(u^{\prime})$. Thus, if we ensure that
$L_{f_{L}}^{(\gamma)}(x)||x-x^{\prime}||_{p}\leq R_{f_{\text{R}}}(f_{\text{L}}(x))$
$\leftrightarrow||x-x^{\prime}||_{p}\leq\frac{R_{f_{R}}(f_{L}(x))}{L_{f_{L}}^{(\gamma)}(x)}$ (11)
then we have:
$g_{SPLITZ}(x)=g_{SPLITZ}(x^{\prime}).$ (12)
However, we also note that:
$||x-x^{\prime}||_{p}\leq\gamma,$ (13)
therefore the certified radius is given by:
$\min\left\{\frac{R_{f_{R}}(f_{L}(x))}{L_{f_{L}}^{(\gamma)}(x)},\gamma\right\}.$ (14)
We finally note that the choice of $\gamma$ (size of the ball) was arbitrary, and we can pick the optimum choice that yields the largest radius. This leads to the final expression for certified radius for SPLITZ:
$R_{g_{\textit{SPLITZ}}}(x)=\underset{\gamma\geq 0}{\max}~{}~{}\min\left\{\frac{R_{f_{\text{R}}}(f_{\text{L}}(x))}{L_{f_{\text{L}}}^{(\gamma)}(x)},\gamma\right\}$ (15)
and completes the proof of the Theorem. ∎
We show the inference and certification process of SPLITZ in Algorithm 1.
Training Methodology for SPLITZ In this Section, we present the details on training the SPLITZ classifier. The key to ensuring the certified robustness of the SPLITZ classifier is to keep the local Lipschitz constant of the left half of the classifier $f_{L}$ small while smoothing the right half of the classifier $f_{R}$. Let us denote $w_{L}$, $w_{R}$ as the training parameters of $f_{L}$ and $f_{R}$, respectively. We propose the following training loss function:
$\underset{w_{L},w_{R}}{\text{min}}~{}\frac{1-\lambda}{N}\sum_{i=1}^{N}\mathbb{E}_{\delta}[\text{Loss}(f_{R}^{w_{R}}(f_{L}^{w_{L}}(x_{i})+\delta),y_{i})]$
$+\frac{\lambda}{N}\sum_{i=1}^{N}\text{max}(\theta,L_{f_{L}^{w_{L}}}^{(\gamma)}(x_{i})),$ (16)
where $\lambda\in[0,1]$ is a hyperparameter controlling the tradeoff between the accuracy and robustness, $\theta$ is a learnable parameter to optimize the local Lipschitz constant, and $\text{Loss}(\cdot)$ is the loss function (e.g., cross entropy loss). Following the literature on randomized smoothing, we replace the expectation operators with their empirical estimates, and our loss function becomes:
$\underset{w_{L},w_{R}}{\text{min}}~{}\underbrace{\frac{1-\lambda}{N}\sum_{i=1}^{N}\left(\frac{1}{Q}\sum_{q=1}^{Q}\text{Loss}(f_{R}^{w_{R}}(f_{L}^{w_{L}}(x_{i})+\delta_{q}),y_{i})\right)}_{\text{"Smoothing" loss for}f_{R}}$
$+\underbrace{\frac{\lambda}{N}\sum_{i=1}^{N}\text{max}(\theta,L_{f_{L}^{w_{L}}}^{(\gamma)}(x_{i}))}_{\text{Lipschitz regularization for}f_{L}}.$ (17)
As illustrated in Eq III, we first input the image $x_{i}$ to the left part of the classifier $f_{L}(x_{i})$ and then add noise $\delta_{q}$, which forms the noisy samples $f_{L}(x_{i})+\delta_{q}$. We then feed noisy samples to the right part of the classifier $f_{R}(f_{L}(x_{i})+\delta_{q})$ and obtain the corresponding prediction. Given the true label $y_{i}$, the loss (e.g., cross entropy) w.r.t $x_{i}$ can be computed. At the same time, the local Lipschitz constant of the left part of the classifier needs to be minimized. To this end, we propose a regularization term to the loss function, which controls the local Lipschitz constant of $f_{L}$. In addition, we do not want the value of the local Lipschitz constant to become too small during the training process, which may lead to a poor accuracy. Therefore, we set a learnable Lipschitz constant threshold $\theta$ for local Lipschitz constant of $f_{L}$, and use $\text{max}(\theta,L_{f_{L}}^{(\gamma)}(x_{i}))$ as the regularization term.
Algorithm 1 SPLITZ Inference & Certification
  Input: Test input $x$, classifier $f^{*}_{L},f^{*}_{R}$, noise level $\sigma$, hyperparameter $\gamma$, confidence parameter $\alpha$, number of noise samples to predict top class $n_{0}$, number of noise samples to estimate the lower bound of the probability of the top class $n_{1}$.
  Output: The certified radius $R_{g_{\textit{SPLITZ}}}(x)$ and corresponding prediction of the given input $x$.
1:  Run the SPLITZ classifier $n_{0}$ times: $\{f^{*}_{R}(f^{*}_{L}(x)+\delta_{i})\}_{i=1}^{n_{0}}$ using $n_{0}$ independent noise realizations.
2:  Compute $\text{count}_{n_{0}}(i)=\text{\# of times class $i$ is the top class}$. SPLITZ inference/prediction: $c_{A}=\arg\max_{i}\text{count}_{n_{0}}(i)$
3:  Approximate the lower confidence bound $\underline{p_{A}}$ of the probability of the top class $c_{A}$ from $n_{1}$ independent runs of the SPLITZ classifier with confidence $1-\alpha$.
4:  if $\underline{p_{A}}>0.5$ then
5:     Compute the certified radius of $f^{*}_{R}$: $R_{f^{*}_{R}}(f^{*}_{L}(x))\leftarrow\sigma\Phi^{-1}(\underline{p_{A}})$
6:     Optimize $\gamma^{*}$ (discussed in Remark $1$) and calculate the corresponding local Lipschitz bound on $f^{*}_{L}$. (Eq. 18).
7:     Compute the overall certified radius $R_{g_{\textit{SPLITZ}}}(x)$ at $x$ (as shown in Theorem 1): $R_{g_{\textit{SPLITZ}}}(x)\leftarrow\min\left\{R_{f^{*}_{\text{R}}}(f^{*}_{\text{L}}(x))/L_{f^{*}_{\text{L}}}^{(\gamma^{*})}(x),\gamma^{*}\right\}$
8:     Return prediction $c_{A}$ and certified radius $R_{g_{\textit{SPLITZ}}}(x)$
9:  else
10:     Return abstain
11:  end if
Computing the Local Lipschitz bound: We note that both SPLITZ training as well as inference/certification require the computation of the local Lipschitz constant of the left half of the network, i.e., $f_{L}$. The simplest approach would be to use a bound on the global Lipschitz constant of $f_{L}$. For example, if $f_{L}$ is composed of $s$ layers, with each layer being a combination of an affine operation followed by ReLU nonlinearity, then the following simple bound could be used:
$L^{(\gamma)}_{f_{L}}(x)\leq||W_{s}||_{2}\times||W_{s-1}||_{2}\dots||W_{1}||_{2},$
where $W_{s}$ is the weight matrix of layer $s$ and $||W_{s}||_{2}$ denotes the corresponding spectral norm. However, this bound, while easy to compute turns out to be quite loose. More importantly, it does not depend on the specific input $x$ as well as the parameter $\gamma$. Fortunately, bounding the local Lipschitz constant of a classifier is an important and a well studied problem. There are plenty of mechanisms to estimate the local Lipschitz bound of $f_{L}$. In principle, our SPLITZ classifier is compatible with these local Lipschitz bound estimation algorithms. From the scope of this paper, we use the local Lischitz constant constrained methodology proposed in [19] which leads to much tighter bounds on the local Lipschitz constant and maintain the specificity on the input $x$. Specifically, we apply the clipped version of activation layers (e.g. ReLU) to constrain each affine layer’s output and obtain the corresponding upper bound (UB) and the lower bound (LB) for each affine layer, where the classifier is given an input $x$ around a $\gamma$ ball. We use an indicator function $I^{v}$ to represent index of the rows or columns in the weight matrices of each affine layer, which within the range from LB to UB. By multiplying each affine layer’s weight matrix and each clipped activation layer’ indicator matrix, the tighter local Lipschtz constant can be obtained. Assume $f_{L}$ network contains $s$-affine-layer neural network and each affine layer is followed by a clipped version of the activation layer, (upper bound of) the local Lipschitz constant $L$ of $f_{L}$ around the input $x$ is:
$L_{f_{L}}^{(\gamma)}(x)\leq$
$\parallel W_{s}I^{v}_{s-1}\parallel_{2}\times\parallel I^{v}_{s-1}W_{s-1}I^{v}_{s-2}\parallel_{2}\dots\parallel I^{v}_{1}W_{1}\parallel_{2},$ (18)
where $W_{s}$ is the weight matrix of layer $s$.
Algorithm 2 SPLITZ Training
  Input: Training set $D_{train}=\{x_{i},y_{i}\}_{i=1}^{N_{train}}$; noise level $\sigma$, training steps $T$, Lipschitz threshold $\theta$, training hyperparameter $\gamma$.
  Output: $f_{L}^{*}$, $f_{R}^{*}$.
1:  for $t=0,\dots,T-1$ do
2:     Compute local Lipschitz constant of $f_{L}$: $L_{f_{L}}^{\gamma}(x)\leftarrow Cal\_Lip~{}(f_{L},x,\gamma)$.
3:     Sample noise $\delta$ and add it to outputs of $f_{L}$ to obtain noise samples: $f_{L}(x)+\delta$
4:     Feed the noise samples to $f_{R}$ network to get the corresponding predictions: $f_{R}(f_{L}(x)+\delta)$
5:     Set the local Lipschitz threshold $\theta$ and minimize the loss function in Eq. III.
6:  end for
  Function $Cal\_Lip$ ($f_{L},x,\gamma$)
1:  Compute the $UB_{k}$ and $LB_{k}$ for each layer $k$ in $f_{L}(x)$ given the perturbation $\gamma$ around input $x$
2:  Compute the indicator matrix $I^{v}_{k}$ for each layer $k$
3:  Compute the local Lipschitz constant $L_{f_{L}}^{(\gamma)}(x)$ (Eq. 18)
  Return $L_{f_{L}}^{(\gamma)}(x)$
Summary of SPLITZ Training Methodology Overall, our training procedure is presented in Algorithm 2. During the process of computing local Lipschitz constant of $f_{L}$, for each iteration, we feed the input to the classifier $f_{L}$ and calculate the LB and UB of outputs of each affine layer in $f_{L}$ given the input $x$ within a $\gamma$ ball. We then can calculate the indicator matrix $I^{v}$ and compute the spectral norm of the reduced weight matrix $\parallel I^{v}_{s}W_{s}I^{v}_{s}\parallel$ for each layer $s$ in $f_{L}$ using power iteration. By multiplying the reduced weight matrix of each affine layer in $f_{L}$, we are able to arrive at the local Lipschitz constant of $f_{L}$. Secondly, we smooth the right half of the neural network $f_{R}$ by sampling from Gaussian noise with zero mean and adding it to the output of $f_{L}$. Then we feed the noisy samples $f_{L}(x)+\delta$ to $f_{R}$ and obtain the corresponding loss. Next, we minimize the overall loss and backward the parameters to optimize the overall network $f$. Finally, we certify the base classifier $f$ to obtain the Lipschitz smoothing classifier $g_{\textit{SPLITZ}}$ as shown in Algorithm 1.
IV Experiments and Evaluation Results
In this section, we evaluate the SPLITZ classifier on two datasets, MNIST[36] and CIFAR-10 [37], where we demonstrate that our proposed approach consistently surpasses other state-of-the-art methods. For all datasets, we report the approximate certified test accuracy and certified radius of smoothed classifiers over test datasets (full test datasets in MNIST and CIFAR-10 datasets). Same as previous works, we vary the noise level $\sigma\in\{0.25,0.5,1.0\}$ for the smoothed models on CIFAR-10 dataset ($\sigma\in\{0.25,0.5,0.75,1.0\}$ for MNIST dataset) and certified the same noise level $\sigma$ during the inference time. To ensure a fair comparison with previous studies, we provide the highest reported results from each paper for the corresponding above levels of noise magnitudes. To improve certified robustness, we utilize the tighter local Lipschitz bound introduced in [19]. For two datasets, we use the same model as previous works [7, 14, 12, 22] (LeNet for MNIST, ResNet110 for CIFAR-10). More experimental details are described in Appendix C.
Evaluation metric Our evaluation metric to measure the certified robustness of the smooth classifier is based on the standard metric proposed in [7]: the approximate certified test accuracy, which can be estimated by the fraction of the test dataset which CERTIFY classifies are correctly classified and at the same time corresponding radius are larger than radius $\epsilon$ without abstaining. Another alternative metric is to measure the average certified radius (ACR) considered by [16]. We show that SPLITZ consistently outperforms other mechanisms w.r.t ACR. For all experiments, we applied the $\ell_{2}$ norm input perturbation.
TABLE II: Comparison of the approximate certified test accuracy (%) on CIFAR-10 under $\ell_{2}$ norm perturbation. Extra data indicates whether their models incorporate other datasets in their models. Each entry lists the certified accuracy using numbers taken from respective papers. We set our values bold face when the value outperforms the best among the comparison methods.
Certified accuracy at $\epsilon~{}(\%)$
Method Extra data 0.25 0.5 0.75 1.0
PixelDP [8] 22.0 2.0 0.0 0.0
RS [7] 61.0 43.0 32.0 22.0
SmoothAdv [15] 67.4 57.6 47.8 38.3
SmoothAdv [15] 74.9 63.4 51.9 39.6
MACER [16] 71.0 59.0 46.0 38.0
Consistency [22] 68.8 58.1 48.5 37.8
SmoothMix [12] 67.9 57.9 47.7 37.2
Boosting [17] 70.6 60.4 52.4 38.8
DRT [23] 70.4 60.2 50.5 39.8
ACES [38] 69.0 57.2 47.0 37.8
DDS [14] 76.7 63.0 45.3 32.1
DDS (finetuning) [14] 79.3 65.5 48.7 35.5
RS + OrthoNN [21] 63.3 45.9 28.9 19.2
SPLITZ 71.3 63.2 53.4 43.2
SPLITZ Methodology For all datasets, we split the classifier after $1^{\text{st}}$ affine layer where the left half of the classifier contains one convolution layer followed by the clipped ReLU layer (See Appendix B). The rest of the classifier is the same as original models (LeNet for MNIST, ResNet110 for CIFAR-10).
Dataset Configuration For both the MNIST and CIFAR-10 dataset(s), we draw $n_{0}=10^{2},~{}n_{1}=10^{5}$ noise samples to certify the smoothing model following [7, 14, 12]. We set the Lipschitz threshold (see Sec III) as $\theta=0.5$. For local Lipschitz constrained training, we set tradeoff parameter $\lambda$ (see Sec III) evenly decrease from $0.8-0.4$, $0.7-0.5$ respectively. We use one Nvidia P100 GPU to train the SPLITZ model with batch size $512,256$ respectively. We apply Adam Optimizer for two datasets (MNIST, CIFAR-10). For the MNIST dataset, we conduct training for 150 epochs and utilize an early-stop strategy to search for the optimal classifier over an additional 150 epochs. We set the initial learning rate as 0.001. The learning rate is decayed (multiplied by 0.1) by 0.1 at every 50 epochs (50th, 100th…). For the CIFAR-10 dataset, we train 200 epochs using the ResNet110 and use the early-stop strategy to search for the optimal classifier over an additional 200 epochs. Furthermore, we set the initial learning rate as 0.001 and final learning rate as $10^{-6}$. The learning rate starts to evenly decay at each epoch from the half of the training epochs. We report our more experimental details in Appendix C.
Baseline mechanisms We compare our method with various existing techniques proposed for robust training of smoothed classifiers, as listed below: (a) PixelDP [8]: cerified robust training with differential privacy mechanism; (b) RS [7]: standard randomized smoothing with the classifier trained with Gaussian augmentation; (c) SmoothAdv [15]: adversarial training combined with randomized smoothing; (d) MACER [16]: a regularization approach which maximizes the approximate certified radius; (e) Consistency [22]: a KL-divergence based regularization that minimizes the variance of smoothed classifiers $f(x+\delta)$ across $\delta$; (f) SmoothMix [12]: training on convex combinations of samples and corresponding adversarial on smoothed classifier; (g) Boosting [17]: a soft-ensemble scheme on smooth training; (h) DRT [23]: a lightweight regularized training on robust ensemble ML models; (i) ACES [38]: a selection-mechanism combined with a smoothed classifier; (j) DDS [14]: a denoised diffusion mechanism combined with a smoothed classifier; (k) RS [7] + OrthoNN [21]: an orthogonal convolutional layer followed by a randomized smoothing model.
IV-A Main Results
Results on MNIST As showed in Table I, we can observe that SPLITZ outperforms other state-of-art approaches in almost every value of $\epsilon$. Impressively, we find that the SPLITZ classifier has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=2.50$, SPLITZ classifier achieves 40.5% compared to state-of-art top-1 test accuracy 34.8% certified test accuracy on the MNIST dataset. Moreover, when we increase $\epsilon$ from 1.50 to 2.50, RS drops from 67.3% to 10.9% decreasing 56.4% test accuracy. SPLITZ, however, maintains higher certified test accuracy from 80.2% to 40.5% maintaining relatively higher test accuracy.
Results on CIFAR10 As shown in Table II and Fig 3, our method outperforms the state-of-art approaches for every value of $\epsilon$ on CIFAR-10 dataset. Interestingly, we find that the SPLITZ training has a significant improvement when the value of $\epsilon$ is large. For instance, when $\epsilon=1.0$, the model achieves 43.2% top-1 test accuracy on CIFAR-10 dataset compared to state-of-art top-1 test accuracy 39.8%. One hypothesis is that minimizing the Lipscitz bound of $f_{L}$ ($L_{f_{L}}\leq 1$) is able to boost the certified radius of the model. Intuitively, more samples are correctly classified while corresponding radius are larger than given $\epsilon$. In addition, we can observe the similar trend as MNIST dataset. SPLITZ maintains higher certified test accuracy when we increase $\epsilon$ from 0.25 to 1.00 compared to other state-of-art mechanisms.
TABLE III: Comparison of average certified radius (ACR) across two different datasets (MNIST, CIFAR-10). We can observe that for two datasets, SPLITZ consistently achieves better results compared to other state-of-art mechanisms. * is reported by [22, 12]
$\sigma$ Methods MNIST CIFAR-10
0.25 $\text{RS}^{*}$ 0.911 0.424
$\text{SmoothAdv}^{*}$ 0.932 0.544
MACER 0.918 0.556
Consistency 0.928 0.552
SmoothMix 0.933 0.548
SPLITZ 1.752 0.680
0.50 $\text{RS}^{*}$ 1.553 0.525
$\text{SmoothAdv}^{*}$ 1.687 0.684
MACER 1.583 0.726
Consistency 1.697 0.726
SmoothMix 1.694 0.737
SPLITZ 2.059 0.924
1.00 $\text{RS}^{*}$ 1.620 0.542
SmoothAdv 1.779 0.660
MACER 1.520 0.792
Consistency 1.819 0.816
SmoothMix 1.823 0.773
SPLITZ 2.104 0.979
Results on Average Certified Radius (ACR) We investigate the performance of SPLITZ using average certified radius (ACR), where we measure the correct samples’ average certified radius over the test datasets (MNIST, CIFAR-10). As shown in Table III, we provide the comprehensive comparison results of average certified radius (ACR) compared to other certified robust techniques. Our SPLITZ consistently outperforms others. For instance, when $\sigma=0.25~{}(0.50)$, the ACR of SPLITZ is 1.752 (2.059) respectively, where the state-of-the-art is 1.697 (0.933) respectively. In addition, we can observe the same trends in CIFAR-10 dataset, where SPLITZ consistently outperforms the state-of-the-art when $\sigma=0.25,0.50$ and $1.00$.
TABLE IV: SPLITZ training time and certifying time across two datasets (MNIST using LeNet, CIFAR-10 using ResNet 110 ).
Training
Datasets Architecture # of GPUs Training per epoch (s) Certifying per image (s)
MNIST LeNet 1 15.1 1.0
CIFAR-10 ResNet110 1 59.5 50.7
TABLE V: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying the splitting layer on MNIST dataset with LeNet.
Certified Test Accuracy at $\epsilon~{}(\%)$
Location of Splitting 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.50
$1^{\text{st}}$ affine layer 94.1 92.0 88.8 84.7 79.0 71.3 62.3 51.7 40.5
$2^{\text{nd}}$ affine layer 89.7 86.2 81.0 75.0 68.0 59.9 51.0 41.4 31.8
$3^{\text{rd}}$ affine layer 84.4 80.4 75.6 69.9 63.6 56.9 49.4 40.3 29.3
IV-B Training and certifying time
Our SPLITZ model needs to vary the value of $\gamma$ (the size of the ball around input $x$) during the training epoch. Thus, we may need relatively more time to obtain the optimal model. To solve this, we apply the early stop mechanism to obtain the better optimized model during the training process. At the same time, we decay our learning rates during the training process. We use one Nvidia P100 GPU to train our SPLITZ model and report our training time (certifying time) for each one epoch (one image) in the Table IV.
IV-C Ablation Study
We also conduct an ablation study to explore the effects of hyperparameters in our proposed method on CIFAR-10 and MNIST datasets. We will explain the impact of the splitting location, the effect of global (local) Lipschitz bound, comparison of the orthogonal neural network (OrthoNN) and SPLITZ, effect of input perturbation $\gamma$ and effect of learnable Lipschitz threshold parameter $\theta$.
Figure 4: (a) Comparison of certified test accuracy of SPLITZ when varying $\sigma$ on the MNIST dataset, (b) Comparison of certified test accuracy of SPLITZ with Global vs Local Lipschitz bound ($\gamma$) on the MNIST dataset.
Impact of splitting location As mentioned in Section III, our SPLITZ classifier can be optimized over different split ways, where we conduct the experiments and show our results in Table V. For example, when $\epsilon=2$, splitting after the $1^{st}$, or $2^{nd}$, or $3^{rd}$ layer result in certified accuracy of $62.3$%, $51.0$% and $49.4$% respectively. These results indicate that splitting the neural network early achieves better performance. Intuitively, splitting the neural network early helps the model minimize the local Lipschitz bound, which improves the certified robustness leading to a higher certified test accuracy given the same $\epsilon$. As the splitting becomes ”deeper”, estimating the local Lipshitz constant also becomes harder, which implies that a looser bound leads to smaller certified radius.
Effect of global (local) Lipschitz constant of the first half of the classifier As shown in Fig 4 (a), we investigate the effect of (upper bound of) the Lipschitz constant of left half of the classifier on certified test accuracy. Interestingly, we can observe that tighter Lipschitz bound gives better certified accuracy given the same radius. Furthermore, using a bound on the local Lipschitz constant to compute the certified accuracy is always better than using the global Lipschitz constant. This is also clearly evident from the result of Theorem 1.
Comparison of the orthogonal neural network (OrthoNN) and the SPLITZ In the baseline RS+OrthoNN model, during the training process, we enforce the first convolution layer of the classifier’s left half to act as an orthogonal convolution neural network. This is achieved by incorporating an orthogonal constraint regularization into the loss function. For the SPLITZ model, we impose a constraint on the local Lipschitz constant of the classifier’s left half. The comparison between RS+orthoNN and SPLITZ is illustrated in Table I with respect to the MNIST dataset and in Table II with respect to the CIFAR-10 dataset. Our observations indicate that enforcing an orthogonal convolution layer on the classifier’s left half has a comparatively lesser impact than training with a local Lipschitz constraint.
TABLE VI: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $Lip_{\theta}$ (the threshold of the local Lipschitz bound around input $x$) on MNIST dataset.
Certified Test Accuracy at $\epsilon~{}(\%)$
$Lip_{\theta}$ 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.50
0.3 94.1 91.6 88.4 84.2 78.2 70.3 60.9 50.6 39.4
0.5 94.1 92.0 88.8 84.7 79.0 71.3 62.3 51.7 40.5
0.7 94.7 92.6 89.4 85.1 79.5 71.5 61.7 50.7 38.8
TABLE VII: Comparison of certified test accuracy of SPLITZ with Gaussian noise $\sigma=0.75$ for varying $\gamma$ (the size of the ball around input $x$) on MNIST dataset.
Certified Test Accuracy at $\epsilon~{}(\%)$
$\gamma$ 0.50 0.75 1.00 1.25 1.50 1.75 2.00 2.25 2.50
0.1 94.7 92.5 89.6 85.4 80.2 73.2 64.0 54.1 42.7
1 94.1 92.0 88.8 84.7 79.0 71.3 62.3 51.7 40.5
2 94.3 92.1 88.6 84.2 78.3 70.4 60.5 49.6 38.0
Effect of $\theta$ (Lipschitz threshold) As shown in Fig 4(b) and Table VI, we analyze the effect of the training threshold $\theta$ (See Eq III). For smaller values of $\epsilon$, SPLITZ with higher Lipschitz constant achieves better performances on accuracy. Conversely, SPLITZ with a smaller Lipschitz constant can boost certified radius, which obtains a relative higher certified test accuracy when $\epsilon$ is larger. These observations suggest a trade-off between the Lipschitz constant and certified accuracy. Thus, identifying the optimal Lipschitz threshold is essential for achieving a balances between robustness and utility. This ablation study further validates that the essence of our SPLITZ classifier lies in maintaining a relatively optimized Lipschitz constant for the left half of the classifier.
Effect of $\gamma$ (size of radius around input $x$) According to above results, constraining the local Lipschitz constant achieves better performance. To further explore the benefit of local Lipschitz constrained training, it is necessary to explore the indicator matrix $I^{v}$ in Eq 18, which depends on the size of the ball around the input (i.e., the hyperparameter $\gamma$). In Table VII, we show how varying training $\gamma$ impacts the certified test accuracy for different values of $\epsilon$. We observe that smaller values of training $\gamma$ lead to higher certified accuracy for all values of $\epsilon$.
V Discussion and conclusion
In this paper, we presented SPLITZ, a novel and practical certified defense mechanism, where we constrained the local Lipschitz bound of the left half of the classifier and smoothed the right half of the classifier with noise. This is because the local Lipschitz constant can capture information specific to each individual input, and the relative stability of the model around that input. To the best of our knowledge, this is the first systematic framework to combine the ideas of local Lipschitz constants with randomized smoothing. Furthermore, we provide a closed-form expression for the certified radius based on the local Lipschitz constant of the left half of the classifier and the randomized smoothing based radius of the right half of the classifier. We show that maintaining a relatively small local Lipschitz constant of the left half of the classifier helps to improve the certified robustness (radius). We showed results on several benchmark datasets and obtained significant improvements over state-of-art methods for MNIST, CIFAR-10 datasets. For instance, on the CIFAR-10 dataset, SPLITZ can achieve $43.2\%$ certified test accuracy compared to state-of-art certified test accuracy $39.8\%$ with $\ell_{2}$ norm perturbation budget of $\epsilon=1$. We reported similar results for the MNIST dataset. We believe that combining the core idea of SPLITZ with other recent techniques, such as denoising diffusion models and adversarial training can be a fruitful next step to further improve certified robustness.
Figure 5: Lipschitz Constants of each affine layer in pretrained models:(a) 4C3F model (there are 4 convolution layers and 3 fully- connected layers in the neural network. ) [19] , (b) Alexnet model [39], (c) ResNet34 model [25]. We can observe the similar trends that right half of the model usually contain a larger Lipschitz constant, while the left half of the model preserves a relatively smaller Lipschitz constant.
Figure 6: (a) Local Lipschitz bound for three random CIFAR-10 images on Alexnet, (b) Percentage analysis of local Lipschitz bound in CIFAR-10 test data.
Appendix A Additional Results on Lipschitz bounds
In this Section, we provide additional Lipschitz constants results in the prevalent neural networks in Fig 5. We can observe the similar trends as previous that the right half of the neural network is more unstable than the right half of the neural network. As shown in Fig 6 (a), we notice considerable variation in the values of local Lipschitz constants across different input images, a trend that is consistent throughout the entire CIFAR-10 test dataset as depicted in Fig 6 (b). These findings lead us to reconsider the efficacy of directly smoothing the input. Such an approach doesn’t cater to the observed heterogeneity. Alternatively, injecting noise at an intermediary step within the classifier can make the model more robust to disturbances.
Appendix B Lipschitz Constrained Training
[…]
$ReLU\theta(x)=\begin{cases}0,&\text{if}~{}~{}x\leq 0\\ x,&\text{if}~{}~{}0<x<\theta\\ \theta,&\text{if}~{}~{}x\geq\theta\\ \end{cases}$ (19)
The neural network maps input $x$ to output $f(x)$ using the following architecture:
$z_{1}=x;z_{l}(x)=ReLU\theta(W_{l}x),z_{L+1}=W_{L}z_{L}$ (20)
We define the perturbation around the input $x$ as:
$x^{\prime}=x+\epsilon,\parallel\epsilon\parallel\leq\delta,\delta\geq 0$ (21)
[…]
$I^{v}(i,i)=\begin{cases}1,&\text{if}~{}~{}UB_{i}>0~{}~{}and~{}~{}LB_{i}<\theta\\ 0,&\text{otherwise}\\ \end{cases}$ (22)
[…]
$D^{v}(i,i)=\begin{cases}\mathbbm{1}(\text{ReLU}\theta(z_{l}^{i})>0),&\text{if}~{}~{}I^{v}(i,i)=1\\ 0,&\text{otherwise}\\ \end{cases}$ (23)
where $\mathbbm{1}$ denote the indicator function. Then the local Lipschitz bound at input $x$ is:
$L_{local}(x,f)\leq\parallel W_{L}I^{v}_{L-1}\parallel\parallel I^{v}_{L-1}W_{L-1}I^{v}_{L-2}\parallel\dots\parallel I^{v}_{1}W_{1}\parallel$ (24)
As stated in [19], it straight forward to prove $\parallel I^{v}_{L-1}W_{L-1}I^{v}_{L-2}\parallel\leq\parallel W_{L-1}\parallel$ using the property of eigenvalues. We briefly prove it following from [19].
Proof.
Let $W^{\prime}=[W~{}I]^{T}$, The singular value of $W^{\prime}$ is defined as the square roots of the eigenvalues of $W^{\prime T}~{}W^{\prime}$. We know the following
${W^{\prime}}^{T}W^{\prime}=W^{T}W+I^{T}I\geq W^{T}W.$ (25)
Therefore, we get the following result:
$\parallel W^{\prime}\parallel\geq\parallel W\parallel$ (26)
We complete the proof. ∎
Next, we will give a toy example to further illustrate the idea of local Lipschitz bound.
[…]
$x\rightarrow\text{Linear1}(W^{1})\rightarrow\text{ReLU}\theta\rightarrow\text{Linear2}(W^{2})\rightarrow y$ (27)
where $x\in\mathcal{R}^{3}$ and $y\in\mathcal{R}$ and $W^{l}$ denotes the weight matrix for layer $l$. Moreover the threshold $\theta=1$.
Given the input [1,-1,0] with $\ell_{2}$ perturbation 0.1. Assume the weight matrices are:
$W^{1}=\begin{bmatrix}2&0&0\\ 0&2&0\\ 0&0&1\end{bmatrix},W^{2}=[1,1,1]$ (28)
Thus, we have the following:
$\text{Input}~{}[1,-1,0]\rightarrow$ (29)
$\begin{bmatrix}[0.9&1.1]\\ [-1.1&-0.9]\\ [-0.1&0.1]\end{bmatrix}\times\begin{bmatrix}2&0&0\\ 0&2&0\\ 0&0&1\end{bmatrix}\rightarrow\begin{bmatrix}[1.8&2.2]\\ [-2.2&-1.8]\\ [-0.1&0.1]\end{bmatrix}$
According to the above upper bound (UB) and lower bound (LB), we obtain the $I_{V}$ function as follows:
$I_{V}^{1}=\begin{bmatrix}0&0&0\\ 0&0&0\\ 0&0&1\end{bmatrix}$ (30)
Overall, we have the local Lipschitz bound as follows:
$L_{local}(x,f)\leq\parallel W^{2}I_{V}^{1}\parallel\parallel I_{V}^{1}W^{1}\parallel=1$ (31)
For the global Lipschitz bound, we have the following:
$L_{global}\leq\parallel W^{2}\parallel\parallel W^{1}\parallel=4$ (32)
Overall, we can find that the local Lipschiz bound is much tighter than the global Lipschitz bound.
Appendix C Additional Experimental Details
In this section, we provide additional results for the two datasets, e.g., MNIST, CIFAR-10 dataset. We first provide the details of two datasets. Next, we illustrate the baselines used in our paper. Note that we report the numbers (certified test accuracy, average certified radius) from respective papers.
C-A Training details
For all value of $\sigma$, we keep the value of training $\sigma$ and testing $\sigma$ to be the same. We apply the noise samples $n_{0}=100$ to predict the most probably class $c_{A}$ and denote $\alpha=0.001$ as the confidence during the certifying process. Furthermore, we use $n_{1}=100,000$ to calculate the lower bound of the probability $p_{A}$ for the MNIST, CIFAR-10 respectively. Moreover, to maintain a relatively small local Lipschitz constant of left half of the SPLITZ classifier, we set the threshold of clipped ReLU (see Sec B) as 1 for all datasets. For estimating the local Lipschitz constant of the left half of the classifier, the power iteration is 5, 2 during the training for MNIST, CIFAR-10 respectively following from [19].
C-B Details of datasets
MNIST [36] dataset contains handwritten digits usually used for image classification problems. The dataset is comprised of a total of 70,000 images, with 60,000 images in the training set and 10,000 in the test set. The dataset has 10 classes and is in grayscale format. We pre-process the MNIST dataset using normalization. We incorporate the Lipschitz constant of the normalization layer when calculating the local Lipschitz constant of the left half of the classifier.
CIFAR-10 [37] dataset consists of 60,000 RGB images distributed across 10 categories: airplane, automobile, bird, cat, deer, dog, frog, horse, ship, and truck. Each category is represented by 6,000 images. The dataset is divided into a training set with 50,000 images and a test set containing 10,000 images. We employ the conventional data augmentation techniques of random horizontal flipping and random translation by up to 4 pixels, consistent with methods used in other baseline studies [14, 7, 12]. Additionally, we normalize each image on a pixel-by-pixel basis. We incorporate the Lipschitz constant of the normalization layer when calculating the local Lipschitz constant of the left half of the classifier.